OpenStudy (anonymous):

Help please... Solve the exponential equation. 1/16 = 64^4x-3

OpenStudy (anonymous):

\[\frac{ 1 }{ 16 }=64^{4x-3}\]

OpenStudy (anonymous):

ok so we better switch to common base 1/16 = 4^(-2) 64^(4x-3) = 4^(3*(4x-3)) can you do it now ?

OpenStudy (anonymous):

So pretty much distribute the 3 to the 4x and -3?

OpenStudy (anonymous):

but then solve for x

OpenStudy (anonymous):

when the base is the same you can equate the the exponents

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

or if you dont want to do it that way u can take the log of both side

OpenStudy (anonymous):

4^(-2) = 4^(3*(4x-3)) 4^(-2) = 4^(12x-9) -2 = 12x-9 7 = 12x x=7/12